OpenStudy (anonymous):

Please Help! Solve the IVP dy/dt - (cos t)/(sin t)*y = 2*sin(t) y(pi/2) = 0 We have been using the Integration factor method

ganeshie8 (ganeshie8):

its a linear ODE, integrting factor will do

ganeshie8 (ganeshie8):

whats your IF here ?

ganeshie8 (ganeshie8):

\[\large \rm y' + \left(- \dfrac{\cos t}{\sin t}\right)y = 2\sin t\]

ganeshie8 (ganeshie8):

\[\large \rm\text{IF} = e^{\int \frac{-cos t}{\sin t} dt} = ?\]

OpenStudy (anonymous):

that's where I get confused it should be e^-csc(t) right...but that makes it super complicated in the end. I don't know what -csc(pi/2) is?

OpenStudy (paxpolaris):

where did you get csc from ?

OpenStudy (paxpolaris):

its not e^-csc(t)

OpenStudy (anonymous):

oh...wait sorry -cot (t)

OpenStudy (anonymous):

ooops....I keep reading it wrong from wolfram

OpenStudy (anonymous):

-cos(t)/sin(t) = -cot(t) and then the integral of -cot(t) is -log(sin(x))

OpenStudy (paxpolaris):

\[\large e^{-\ln(\sin x)}=?\]

OpenStudy (anonymous):

-sin(x) ?

OpenStudy (paxpolaris):

\[\large=\left[ e^{\ln(\sin x)} \right]^{-1}\]

OpenStudy (paxpolaris):

=?

OpenStudy (anonymous):

(sin x)^-1 = 1/sin(x)

OpenStudy (paxpolaris):

right that's your integrating factor ... you multiply to the original equatiom

OpenStudy (anonymous):

ok awesome! I think im right from here thanks! I just got caught up in the integral and trig identities. Need to practise that :)